# Robust Control Barrier Functions using Uncertainty Estimation with Application to Mobile Robots

### v2

@@ Abstract
This paper proposes a safety-critical control design approach for nonlinear control affine systems in the presence of matched and unmatched uncertainties. Our constructive framework couples control barrier function (CBF) theory with a new uncertainty estimator to ensure robust safety. The estimated uncertainty with a derived upper bound on the estimation error is used for synthesizing CBFs and safety-critical controllers via a quadratic program-based feedback control law that rigorously ensures robust safety while improving disturbance rejection performance. The method is extended to higher-order CBFs (HOCBFs) to achieve safety under unmatched uncertainty, which may cause relative degree differences with respect to control input and disturbances. We assume the relative degree difference is at most one, resulting in a second-order cone constraint. The proposed robust HOCBF method is demonstrated via a simulation of an uncertain elastic actuator control problem. Finally, we experimentally demonstrated the efficacy of our robust CBF framework on a tracked robot with slope-induced matched and unmatched perturbations.
I Introduction
@@
We organize this paper as follows. Section II provides preliminaries. Section III proposes and analyzes an uncertainty estimator and introduces estimator-based robust CBF schemes. Section IV presents Simulations and experiments, while Section V concludes the paper.
II Preliminaries
@@
Hence, given any initial condition ${x_{0}\!\triangleq\!x(t_{0})\!\in\!X}$ there exists an interval ${\mathcal{I}({x}_{0})\!\triangleq\!\left[t_{0},t_{\max}\right)}$ such that the solution of (2) produces a unique trajectory for ${t\!\in\!\mathcal{I}({x}_{0})}$. Throughout this study we assume ${f_{\rm cl}}$ is forward complete, i.e., ${\mathcal{I}({x}_{0})\!=\![0,\infty)}$, and ${U}$ is a convex polytope.
II-A Control Barrier Functions
@@
$L_{f}^{i}h(x)\!\triangleq\!\mathcal{J}_{x}L_{f}^{i-1}h(x)f(x),\ L_{g}L_{f}^{i-1}h(x)\!\triangleq\!\mathcal{J}_{x}L_{f}^{i-1}h(x)g(x),$
with ${\mathcal{J}_{x}L_{f}^{i-1}h(x)\!\triangleq\!\frac{\partial L_{f}^{i-1}h(x)}{\partial x}}$, ${L_{f}^{0}h(x)\!\!\triangleq\!h(x)}$, ${L_{f}^{i}h\!:\!X\!\!\to\!\!\mathbb{R}}$, ${\mathcal{J}_{x}L_{f}^{i-1}h\!:\!X\!\!\to\!\!\mathbb{R}^{1\times n}}$, ${L_{g}L_{f}^{i-1}h\!:\!X\!\!\to\!\!\mathbb{R}^{m}}$.
@@
where ${\alpha_{i}\!\in\!\mathcal{K}_{\infty,e}}$ is a ${(r\!-\!i)^{th}}$ order differentiable function. The associated extended sets and their intersections are defined as:
@@
II-B Analysis of the Matched and Unmatched Uncertainties
In practice, uncertainties and disturbances cannot be fully modelled. Thus, functions $f$ and $g$ in the true model (1) may be imprecisely known, leading to degraded system performance and stability. In safe control synthesis, we typically use a control affine nominal model to represent our best understanding of the system:
@@ II-B Analysis of the Matched and Unmatched Uncertainties
The matrix $\hat{g}$ has full column rank, i.e, ${\operatorname{rank}({\hat{g}(x))}\!=\!m,\ \forall x\!\in\!X}$, and the left pseudo-inverse of $\hat{g}$: ${\hat{g}^{\dagger}\!(x)\!=\!\big{(}\hat{g}^{\top}\!(x)\hat{g}(x)\big{)}^{-1}\hat{g}^{\top}(x)}$, with ${\hat{g}^{\dagger}\!:\!X\!\to\!\mathbb{R}^{m\times n}}$, exists ${\forall x\!\in\!X}$.
@@
where ${{f}(x)\!-\!\hat{f}(x)\!\!\triangleq\!\Delta f(x)}$, ${{g}(x)\!-\!\hat{g}(x)\!\!\triangleq\!\Delta g(x)}$; ${\Delta f,\Delta g\!:\!X\!\!\to\!\mathbb{R}^{n}}$, are the unmodelled parts of ${\Delta\!:\!X\!\times\!U\!\to\!\mathbb{R}^{n}}$. Note that model description (11) implicitly ignores time-varying system disturbances, which implies the system (1) is time-invariant. Purely time-varying input disturbances in (9) of the form: ${\dot{x}\!=\!\hat{f}(x)\!+\!\hat{g}(x)u\!+\!d(t),d\!:\!\mathbb{R}^{+}_{0}\!\to\!\mathbb{R}^{n}}$ are assumed locally Lipschitz continuous in $t$ over ${t\geq 0}$.
@@
Assumption 2 is reasonable for many practical nonlinear systems, such as wheeled mobile robots [25], aerial robots [27], and ships [28]. The upper bounds ${\delta_{b},\delta_{L}}$ can be learned from data obtained from system simulations or tests. Moreover, note that ${\delta_{b}}$ need not be known for our safety guarantees with CBFs in our proposed methods, but is needed to characterize the estimation error dynamics.
@@
The disturbance relative degree of a sufficiently differentiable output function ${h\!:\!X\!\to\!\mathbb{R}}$ of the uncertain system (9) on a set ${\mathcal{S}\!\subseteq\!X}$ with respect to $\Delta$, is defined as an integer ${v\!\leq\!n}$ if ${L_{\Delta}L_{\hat{f}}^{v-1}h(t,\!x,\!u)\!\neq\!0}$ for all ${(t,\!x,\!u)\!\in\!\mathbb{R}^{+}_{0}\!\times\!\mathcal{S}\!\times\!U}$, and ${L_{\Delta}L_{\hat{f}}^{i-1}h(t,\!x,\!u)\!=\!0}$, for ${i\!\in\!\{1,\ldots,v-1\}}$, where ${L_{\Delta}L_{\hat{f}}^{i-1}h(t,\!x,\!u)\!\triangleq\!\mathcal{J}_{x}L_{\hat{f}}^{i-1}h(x)\Delta(t,\!x,\!u)}$.
@@
The compound uncertainly $\Delta$ in system (9) satisfies ${\text{DRD}\!\geq\!r\!-\!1}$ when ${\text{IRD}\!=\!r}$ for ${r\!\geq\!2}$ with respect a sufficiently differentiable output function ${h\!:\!X\!\to\!\mathbb{R}}$ of system (9).
@@
$+\dfrac{d}{dt}\!\big{(}\!\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\!\big{)}\!{\Delta}(t,\!x,\!u)\!+\!\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\dot{\Delta}(t,\!x,\!u).$ (17)
One can observe from (II-B) that, if ${\text{IRD}\!=\!\text{DRD}\!=\!r}$, then
@@
which shows that disturbance $\Delta$ only affects the highest-order derivative of $h$. When ${(\text{IRD}\!=\!r)<\text{DRD}}$, $\Delta$ disappears from (II-B): it does not affect the HOCBF conditions.
Ignoring the effects of $\Delta$ in (16) or (II-B) may lead to unsafe behavior. To remedy this problem, we estimate and bound the uncertainty in a quantifiable way, compensate the full-state matched part of the uncertainty, and incorporate the error bound and unmatched part into the CBF and HOCBF conditions. The problem statement of this work:
@@
Given the nominal dynamics ${\dot{x}\!=\!\hat{f}(x)\!+\!\hat{g}(x)u}$ in (9) and the upper bounds ${\delta_{L},\delta_{b}\!\in\!\mathbb{R}^{+}_{0}}$ for ${\Delta}$, which is a composition of output-matched and output-unmatched uncertainties, and a safe set $\mathcal{C}$ defined by (3), synthesize a safe controller under Assumption 3.
III Main Result
@@ III-A Uncertainty Estimator
where ${\hat{\Delta}\!:\!X\!\times\!\mathbb{R}^{n}\!\to\!\mathbb{R}^{n}}$ is the estimated uncertainty, ${\xi\!\in\!\mathbb{R}^{n}}$ is an auxiliary state vector, ${0\!\prec\!\Lambda\!\in\!\mathbb{R}^{n\times n}}$ is a diagonal positive definite design matrix, and ${s_{n}\!\in\!\mathbb{R}^{n}}$ is bounded measurement noise: ${\|s_{n}(t)\|_{\infty}\!\triangleq\!\sup_{t}{\|s_{n}(t)\|}\!\leq\!\bar{s}_{n}}$ for some known ${\bar{s}_{n}\!\in\!\mathbb{R}^{+}_{0}}$. Without loss of generality, we set the initial value of ${\hat{\Delta}}$, ${\hat{\Delta}(0)\!=\!{\bf 0}}$, by assigning ${\xi(0)\!=\!\Lambda x(0)}$ in (18). The uncertainty estimation error is
@@
Consider the uncertain system (9) with a compound uncertainty function ${\Delta\!:\!\mathbb{R}^{+}_{0}\!\times\!X\!\times\!U\!\to\!\mathbb{R}^{n}}$ that satisfies Assumption 2 with upper bounds ${\delta_{L},\delta_{b}\!\in\!\mathbb{R}^{+}_{0}}$, and the uncertainty estimator (18)-(19) with a diagonal positive definite matrix ${\Lambda\!\in\!\mathbb{R}^{n\times n}}$. Then, error $e$ given in (20) is uniformly ultimately bounded.
@@
III-B Uncertainty Estimation-Based Safety with CBFs
To guarantee CBF robustness, we must consider uncertainty ${\Delta}$ in our safety analysis, even though it cannot be measured in practice. To address this issue, we replace ${\Delta}$ with the estimated uncertainty ${\hat{\Delta}}$ and a lower bound on the estimation error ${{e}}$. Theorem 3 summarizes the robust safety assurances provided by uncertainty estimation and full-state matched uncertainty cancellation. Note that here we consider CBFs with ${\text{IRD}\!=\!1}$; therefore, from Assumption 3, ${\text{DRD}\!\geq\!1}$.
@@ III-B Uncertainty Estimation-Based Safety with CBFs
We remark that given a locally Lipschitz continuous nominal controller, $\mathbf{k_{d}}$, Theorem 3 allows for the design of a safe controller by solving an uncertainty estimation and cancellation-based robust safety filter, where the constraint is expressed by (III-B). This approach enables the implementation of the controller as an efficient QP.
Note that the robust CBF condition (III-B) explicitly depends on the time-varying upper error bound, ${\bar{e}}$. A similar method was used in [17] to address robustness against unmodeled system dynamics by estimating the effect of the disturbance on the CBF. While safety was guaranteed in [17], uncertainty compensation via estimation was not considered, and the proposed safety filter may perform poorly.
@@
Condition ${2\sigma_{V}\mu_{e}\!-\!{\sigma_{V}\alpha_{h}}/{2}\!>\!0\!\!\iff\!\!\lambda_{\rm min}(\Lambda)\!>\!\alpha_{h}}$ requires the uncertainty estimation dynamics represented by $\lambda_{\rm min}(\Lambda)$ to be faster than the exponential rate of the CBF constraint, which is characterized by $\alpha_{h}$. For a given uncertainty estimator, it is possible to select an $\alpha_{h}$ value that is sufficiently small in practical applications.
@@
Safety condition (32) is not explicitly affected by the time-varying estimation error $e$, whereas (27) depends on it. Hence, Theorem 4 gives design flexibility even when upper bound ${\delta_{b}}$ is unknown, as the upper bound of $\|e\|$, ${\bar{e}(t)}$ explicitly depends on ${\delta_{b}}$.
@@
III-C Uncertainty Estimation-Based Safety with HOCBFs
@@
Suppose ${\text{IRD}\!=\!r}$ and ${\text{DRD}\!=\!r\!-\!1}$, and ${r\!\geq\!2}$, the higher-order time derivatives of $h$ with respect to system (9) are given in (II-B). We modify (6) with respect to the uncertain system (9), which results in:
@@
where ${\Psi(x)\!\triangleq\!L_{\hat{f}}^{r-1}h(x)\!+\!\sum_{i=1}^{r-1}L_{\hat{f}}^{i-1}(\alpha_{r-i}\circ{\phi}_{r-i-1})(x)}$. Note that, $\Delta$ first appears in ${\phi_{r-1}}$ as it depends on ${{h}^{r-1}}$, which depends on $\Delta$; see (II-B). We also define the following sets and their intersections:
@@
$L_{\hat{f}}\Psi(x)\!+\!L_{\hat{g}}\Psi(x)u\!+\!\nabla\Psi(x){\Delta}(t,\!x,\!u)\!+\!\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\!~{}\dot{\Delta}(t,\!x,\!u)$
@@
$\!\geq\!-\alpha_{r}\big{(}\Psi(x)+\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x){\Delta}(t,\!x,\!u)\big{)},\ \forall x\in\mathcal{C},$ (35)
where ${\alpha_{r}\!\in\!\mathcal{K}_{\infty,e}}$, ${\mathcal{F}(x)\!\triangleq\!\frac{\partial}{\partial x}\big{(}\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\big{)}}$, with ${\mathcal{F}\!:\!X\!\to\!\mathbb{R}^{n\times n}}$. And, from Assumption 2 we have
$\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\!~{}\dot{\Delta}(t,\!x,\!u)\!\geq\!-\big{\|}\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\big{\|}\delta_{L},\forall(t,\!x,\!u)\!\in\!\mathbb{R}^{+}_{0}\!\times\!X\!\times\!U.$
Like Theorem 3, we replace $\Delta$ with ${\hat{\Delta}\!+\!e}$, and $u$ with ${\bar{u}\!-\!\hat{g}^{\dagger}\hat{\Delta}}$ in (III-C) to integrate the uncertainty estimator into the robust safety constraint. This process results in the following error-dependent HOCBF condition similar to (III-B):
$\begin{split}&L_{\hat{f}}\Psi(x)+L_{\hat{g}}\Psi(x)\big{(}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\nabla\Psi(x)\hat{\Delta}(t)\\ &+\big{(}\hat{f}(x)+\hat{g}(x)\big{(}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\hat{\Delta}(t)\big{)}^{\top}\mathcal{F}(x)\hat{\Delta}(t)\\ &\!-\!\bar{e}(t)\big{\|}\big{(}\hat{f}(x)+\hat{g}(x)\big{(}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\hat{\Delta}(t)\big{)}^{\top}\mathcal{F}(x)\big{\|}\\ &-\big{\|}\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\big{\|}\delta_{L}-\Omega(t,x,e)\\ &\geq\!-\alpha_{r}\big{(}\Psi(x)\!+\!\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\hat{\Delta}(t)\!-\!\big{\|}\mathcal{J}_{x}L_{\hat{f}}^{r-2}h(x)\big{\|}\bar{e}(t)\big{)},\end{split}$ (36)

### v3

@@ Abstract
This paper proposes a safety-critical control design approach for nonlinear control affine systems in the presence of matched and unmatched uncertainties. Our constructive framework couples control barrier function (CBF) theory with a new uncertainty estimator to ensure robust safety. We use the estimated uncertainty, along with a derived upper bound on the estimation error, for synthesizing CBFs and safety-critical controllers via a quadratic program-based feedback control law that rigorously ensures robust safety while improving disturbance rejection performance. We extend the method to higher-order CBFs (HOCBFs) to achieve safety under unmatched uncertainty, which may cause relative degree differences with respect to control input and disturbances. We assume the relative degree difference is at most one, resulting in a second-order cone constraint. We demonstrate the proposed robust HOCBF method through a simulation of an uncertain elastic actuator control problem and experimentally validate the efficacy of our robust CBF framework on a tracked robot with slope-induced matched and unmatched perturbations.
{IEEEkeywords}
Control barrier functions, constrained control, robotics, robust control, uncertainty estimation
1 Introduction
@@
We organize this paper as follows. Section 2 provides preliminaries. Section 3 proposes and analyzes an uncertainty estimator and introduces estimator-based robust CBF schemes. Section 4 presents simulations and experiments, while Section 5 concludes the paper.
2 Preliminaries
@@
Hence, given any initial condition ${x_{0}\!\triangleq\!x(t_{0})\!\in\!X}$ there exists an interval ${\mathcal{I}({x}_{0})\!\triangleq\!\left[t_{0},t_{\max}\right)}$ such that the solution of (2) produces a unique trajectory for ${t\!\in\!\mathcal{I}({x}_{0})}$. Throughout this study, we assume ${f_{\rm cl}}$ is forward complete, i.e., ${\mathcal{I}({x}_{0})\!=\![0,\infty)}$, and ${U}$ is a convex polytope.
2.1 Control Barrier Functions
@@
$L_{f}^{i}h(x)\!\triangleq\!\nabla L_{f}^{i-1}h(x)f(x),\ L_{g}L_{f}^{i-1}h(x)\!\triangleq\!\nabla L_{f}^{i-1}h(x)g(x),$
where ${\nabla L_{f}^{i-1}h(x)\!\triangleq\!\frac{\partial L_{f}^{i-1}h(x)}{\partial x}}$, ${L_{f}^{0}h(x)\!\!\triangleq\!h(x)}$, with ${L_{f}^{i}h\!:\!X\!\!\to\!\!\mathbb{R}}$, ${\nabla L_{f}^{i-1}h\!:\!X\!\!\to\!\!\mathbb{R}^{1\times n}}$, and ${L_{g}L_{f}^{i-1}h\!:\!X\!\!\to\!\!\mathbb{R}^{m}}$.
@@
where ${\alpha_{i}\!\in\!\mathcal{K}_{\infty,e}}$ is a ${(r\!-\!i)^{th}}$ order differentiable function. The associated extended sets and their intersections are defined as
@@
2.2 Analysis of the Matched and Unmatched Uncertainties
In practice, uncertainties and disturbances cannot be fully modeled. Thus, functions $f$ and $g$ in the true model (1) may be imprecisely known, leading to degraded system performance and stability. In safe control synthesis, we typically use a control affine nominal model to represent our best understanding of the system:
@@ 2.2 Analysis of the Matched and Unmatched Uncertainties
The matrix $\hat{g}$ has full column rank, i.e., for all ${x\!\in\!X}$: ${\operatorname{rank}({\hat{g}(x))}\!=\!m}$, and the left pseudo-inverse of $\hat{g}$: ${\hat{g}^{\dagger}\!(x)\!=\!\big{(}\hat{g}^{\top}\!(x)\hat{g}(x)\big{)}^{-1}\hat{g}^{\top}(x)}$, with ${\hat{g}^{\dagger}\!:\!X\!\to\!\mathbb{R}^{m\times n}}$, exists.
@@
where ${{f}(x)\!-\!\hat{f}(x)\!\!\triangleq\!\Delta f(x)}$, ${{g}(x)\!-\!\hat{g}(x)\!\!\triangleq\!\Delta g(x)}$ are the unmodelled parts of ${\Delta}$. Note that model description (11) implicitly ignores time-varying system disturbances, which implies the system (1) is time-invariant. Purely time-varying input disturbances in (9) of the form: ${\dot{x}\!=\!\hat{f}(x)\!+\!\hat{g}(x)u\!+\!d(t)}$, with ${d\!:\!\mathbb{R}^{+}_{0}\!\to\!\mathbb{R}^{n}}$, are assumed locally Lipschitz continuous in $t$ over ${t\geq 0}$.
@@
Assumption 2 is reasonable for many practical nonlinear systems, such as wheeled mobile robots [25], aerial robots [27], and ships [28]. The upper bounds ${\delta_{b},\delta_{L}}$ can be learned from data obtained from system simulations or tests. Moreover, note that ${\delta_{b}}$ need not be known for our proposed CBF-based safety guarantees, but is needed to characterize the estimation error dynamics.
@@
The disturbance relative degree of a sufficiently differentiable output function ${h(x),h\!:\!X\!\to\!\mathbb{R},}$ of the uncertain system (9) on a set ${\mathcal{S}\!\subseteq\!X}$ with respect to $\Delta$, is defined as an integer ${v\!\leq\!n}$ if ${L_{\Delta}L_{\hat{f}}^{v-1}h(t,x,u)\!\neq\!0}$, ${\forall(t,x,u)\!\in\!\mathbb{R}^{+}_{0}\!\times\!\mathcal{S}\!\times\!U}$, and ${L_{\Delta}L_{\hat{f}}^{i-1}h(t,x,u)\!\equiv\!0}$, where ${L_{\Delta}L_{\hat{f}}^{i-1}h(t,x,u)\triangleq\nabla L_{\hat{f}}^{i-1}h(x)\Delta}$, for ${i\!\in\!\{1,\ldots,v-1\}}$.
@@
The compound uncertainty $\Delta$ in system (9) satisfies ${\text{DRD}\!\geq\!r\!-\!1}$ when ${\text{IRD}\!=\!r}$ for ${r\!\geq\!2}$ with respect to a sufficiently differentiable output function ${h\!:\!X\!\to\!\mathbb{R}}$ of system (9).
@@
$+\dfrac{d}{dt}\!\big{(}\!\nabla L_{\hat{f}}^{r-2}h(x)\!\big{)}\!{\Delta}(t,\!x,\!u)\!+\!\nabla L_{\hat{f}}^{r-2}h(x)\dot{\Delta}(t,\!x,\!u).$ (17)
One can observe from (2.2) that, if ${\text{IRD}\!=\!\text{DRD}\!=\!r}$, then
@@
which shows that the uncertainty $\Delta$ only affects the highest-order derivative of $h$. When ${(\text{IRD}\!=\!r)<\text{DRD}}$, the term $\Delta$ disappears from (2.2): it does not affect the HOCBF conditions.
Ignoring the effects of $\Delta$ in (16) or (2.2) may lead to unsafe behavior. To remedy this problem, we estimate and bound the uncertainty in a quantifiable way, compensate the full-state matched uncertainty, and incorporate the error bound and unmatched part into the CBF and HOCBF conditions. The problem statement of this work follows:
@@
Given the nominal system dynamics ${\dot{x}\!=\!\hat{f}(x)\!+\!\hat{g}(x)u}$ in (9) and the upper bounds ${\delta_{L},\delta_{b}\!\in\!\mathbb{R}^{+}_{0}}$ for ${\Delta}$, which can be a composition of output-matched and output-unmatched uncertainties with respect to CBF ${h}$, and a safe set $\mathcal{C}$ defined by (3), synthesize a safe controller under Assumption 3.
3 Main Result
@@ 3.1 Uncertainty Estimator
where ${\hat{\Delta}\!:\!X\!\times\!\mathbb{R}^{n}\!\to\!\mathbb{R}^{n}}$ is the estimated uncertainty, ${\xi\!\in\!\mathbb{R}^{n}}$ is an auxiliary state vector, ${0\!\prec\!\Lambda\!\in\!\mathbb{R}^{n\times n}}$ is a positive definite design matrix, and ${s_{n}\!\in\!\mathbb{R}^{n}}$ is bounded measurement noise: ${\|s_{n}(t)\|_{\infty}\!\triangleq\!\sup_{t}{\|s_{n}(t)\|}\!\leq\!\bar{s}_{n}}$ for some known ${\bar{s}_{n}\!\in\!\mathbb{R}^{+}_{0}}$. Without loss of generality, we set the initial value of ${\hat{\Delta}}$, ${\hat{\Delta}(0)\!=\!{\bf 0}}$, by assigning ${\xi(0)\!=\!\Lambda x(0)}$ in (18). The uncertainty estimation error is
@@
Consider the uncertain system (9) with a compound uncertainty function ${\Delta\!:\!\mathbb{R}^{+}_{0}\!\times\!X\!\times\!U\!\to\!\mathbb{R}^{n}}$ that satisfies Assumption 2 with upper bounds ${\delta_{L},\delta_{b}\!\in\!\mathbb{R}^{+}_{0}}$, and the uncertainty estimator (18)-(19) with a positive definite matrix ${\Lambda\!\in\!\mathbb{R}^{n\times n}}$. Then, error $e$ given in (20) is uniformly ultimately bounded.
@@
We can also simply take ${D=1}$ and ${\tau_{e}=\lambda_{\rm min}(\Lambda)}$ if ${\Lambda}$ is symmetric positive definite. ∎
@@
3.2 Uncertainty Estimation-Based Safety with CBFs
To guarantee CBF robustness, we must consider the uncertainty ${\Delta}$ in our safety analysis, even though it cannot be measured in practice. To address this issue, we replace ${\Delta}$ with the estimated uncertainty ${\hat{\Delta}}$ and a lower bound on the estimation error ${{e}}$. Theorem 3 summarizes the robust safety assurances provided by uncertainty estimation and full-state matched uncertainty cancellation. Note that here we consider CBFs with ${\text{IRD}\!=\!1}$; therefore, from Assumption 3, ${\text{DRD}\!\geq\!1}$.
@@ 3.2 Uncertainty Estimation-Based Safety with CBFs
We remark that given a locally Lipschitz continuous nominal controller, $\mathbf{k_{d}}$, Theorem 3 allows for the design of a safe controller by formulating an uncertainty estimation and cancellation-based robust safety filter, where the constraint is expressed by (3.2). This approach enables the implementation of the controller as an efficient QP.
Note that the robust CBF condition (3.2) explicitly depends on the time-varying upper error bound, ${\bar{e}}$. A similar method was used in [17] to address robustness against unmodeled system dynamics by estimating the effect of the disturbance on the CBF. While safety was guaranteed in [17], uncertainty compensation via estimation was not considered, and the proposed safety filter may perform poorly.
@@
Condition ${2\sigma_{V}\mu_{e}\!-\!{\sigma_{V}\alpha_{h}}/{2}\!>\!0\!\!\iff\!\!\lambda_{\rm min}(\Lambda)\!>\!\alpha_{h}}$ requires the uncertainty estimation dynamics represented by $\lambda_{\rm min}(\Lambda)$ to be faster than the exponential rate of the CBF constraint, which is characterized by $\alpha_{h}$. For a given uncertainty estimator, one can select an $\alpha_{h}$ value that is sufficiently small in practical applications.
@@
Safety condition (32) is not explicitly affected by the time-varying estimation error $e$, whereas (27) depends on it. Hence, Theorem 4 gives design flexibility even when the upper bound ${\delta_{b}}$ is unknown, as the upper bound of $\|e\|$, ${\bar{e}(t)}$, explicitly depends on ${\delta_{b}}$.
@@
3.3 Uncertainty Estimation-Based Safety with HOCBFs
@@
Suppose ${\text{IRD}\!=\!r}$ and ${\text{DRD}\!=\!r\!-\!1}$, and ${r\!\geq\!2}$, the higher-order time derivatives of $h$ with respect to system (9) are given in (2.2). We modify (6) with respect to the uncertain system (9), which results in:
@@
where ${\Psi(x)\!\triangleq\!L_{\hat{f}}^{r-1}h(x)\!+\!\sum_{i=1}^{r-1}L_{\hat{f}}^{i-1}(\alpha_{r-i}\circ{\phi}_{r-i-1})(x)}$. Note that $\Delta$ first appears in ${\phi_{r-1}}$ as it depends on ${{h}^{r-1}}$, which depends on $\Delta$; see (2.2). We also define the following sets and their intersections:
@@
$L_{\hat{f}}\Psi(x)\!+\!L_{\hat{g}}\Psi(x)u\!+\!\nabla\Psi(x){\Delta}(t,\!x,\!u)\!+\!\nabla L_{\hat{f}}^{r-2}h(x)\!~{}\dot{\Delta}(t,\!x,\!u)$
@@
$\!\geq\!-\alpha_{r}\big{(}\Psi(x)+\nabla L_{\hat{f}}^{r-2}h(x){\Delta}(t,\!x,\!u)\big{)},\ \forall x\in\mathcal{C},$ (35)
where ${\alpha_{r}\!\in\!\mathcal{K}_{\infty,e}}$, ${\mathcal{F}(x)\!\triangleq\!\frac{\partial}{\partial x}\big{(}\nabla L_{\hat{f}}^{r-2}h(x)\big{)}}$, with ${\mathcal{F}\!:\!X\!\to\!\mathbb{R}^{n\times n}}$. And, from Assumption 2 we have
$\nabla L_{\hat{f}}^{r-2}h(x)\!~{}\dot{\Delta}(t,\!x,\!u)\!\geq\!-\big{\|}\nabla L_{\hat{f}}^{r-2}h(x)\big{\|}\delta_{L},\forall(\!t,\!x,\!u\!)\!\in\!\mathbb{R}^{+}_{0}\!\times\!X\!\times\!U.$
Like Theorem 3, we replace $\Delta$ with ${\hat{\Delta}\!+\!e}$, and $u$ with ${\bar{u}\!-\!\hat{g}^{\dagger}\hat{\Delta}}$ in (3.3) to integrate the uncertainty estimator into the robust safety constraint. This process results in the following error-dependent HOCBF condition similar to (3.2):
$\begin{split}&L_{\hat{f}}\Psi(x)+L_{\hat{g}}\Psi(x)\big{(}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\nabla\Psi(x)\hat{\Delta}(t)\\ &+\big{(}\hat{f}(x)+\hat{g}(x)\big{(}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\hat{\Delta}(t)\big{)}^{\top}\mathcal{F}(x)\hat{\Delta}(t)\\ &\!-\!\bar{e}(t)\big{\|}\big{(}\hat{f}(x)+\hat{g}(x)\big{(}\bar{u}-\hat{g}^{\dagger}\!(x)\hat{\Delta}(t)\big{)}+\hat{\Delta}(t)\big{)}^{\top}\mathcal{F}(x)\big{\|}\\ &-\big{\|}\nabla L_{\hat{f}}^{r-2}h(x)\big{\|}\delta_{L}-\Omega(t,x,e)\\ &\geq\!-\alpha_{r}\big{(}\Psi(x)\!+\!\nabla L_{\hat{f}}^{r-2}h(x)\hat{\Delta}(t)\!-\!\big{\|}\nabla L_{\hat{f}}^{r-2}h(x)\big{\|}\bar{e}(t)\big{)},\end{split}$ (36)